A car with two people in it ( mass 1000kg) is parked 10.0m from edge of cliff. A person ties a rope to the car's bumper and a 50 kg rock to the other end of the rope. He lowers the rock over the edge of the cliff and the car which is in neutral accelerates toward the edge. Determine the acceleration of the car towards the edge and how long to the people in the car have to jump out before the car goes over the edge?
Answer #1

for rope

m*g - T = m*a

T = mg - ma

for car


T = M*a


mg - ma = M*a


(50*9.8) - (50*a) = 1000*a


acceleration = a = 0.47 m/s^2


distance = x = 10 m

x = ut + 0.5*a*t^2

10 = 0 + (0.5*0.47*t^2


t = 6.53 s